OpenStudy (anonymous):

The x-coordinate of the point of intersection of the straight lines 2x+ay-7=0 and ax-y-5=0 is 2. If a is an integer, find the value of a.

OpenStudy (anonymous):

2x+ay-7=0 -(1) ax-y-5=0 -(2) From the second equation: y = ax-5 -(3) Sub. (3) into (1) 2x+a(ax-5)-7=0 -(4) Sub x=2 into (4), then solve a I think it should work.

OpenStudy (anonymous):

is it 3?

OpenStudy (anonymous):

thank you I got it:)

OpenStudy (anonymous):

I think so!